An astronaut with a mass of 70 kg is standing on the surface of a spherical asteroid that is 10 km in diameter and has a mass of 1x1015 kg. He knows he can jump straight up 0.2 meters on Earth with his spacesuit on. Can he generate enough energy by jumping to propel himself high enough off of the surface of the asteroid to grab a spacecraft that is orbiting 1 km above? (G=6.67x10-11Nm2/kg2 , g=9.8m/s2)
Answer #1

The formula for gravitational field due to asteroid at its surface

g_{asteroid} = \frac{GM}{r^{2}}

putting all the values we get

g_{asteroid} = 0.2668*10^{-2} m/s^{2}

maximum height is given as

H_{max}= v^{2}/2a

Keeping v same for earth and asteroid we can see that

H_{asteroid}* g_{asteroid}= H_{earth}*g_{earth}

H_{asteroid}= 0.2*9.8*10^{2} / 0.2668 =0.735 km

no maximum height comes less than 1 km, so its not possible to jump at that height